Example Suppose the buyer normally receives shipments from the four manufacturers in quantities of 10 If all shipments are made less-than-t Proposal. The distribution warehouse suggests consolidation ser 8, 15 and 7 tons. uckload drectly to the customer, the transpotaion aes are 20, 18, 15, 1nd 18 LVL per unit. service and transportation to the customer by one truck. You are the manufacturer. Should you accept such proposal? The rates are manufacture LT L rate to Warehouse FTL rate from wh to warehouse charge customer EUR/TOn EUR EUR/ton 10 10 6,0 12.0 5,0 10 10 10 15 DIBOSS
Answer #1

Since we send four shipments of quantities 10, 8, 15 and 7 tons to the customer, we are also required to bear the transportation expense. When we send four separate truck loads of the shipments, the 4 trucks charges us separately and we are needed to bear that expense. Let us suppose that the total expense on only transporting the shipments via the 4 trucks is (x+x+x+x) = 4x.

                  Now, if we apply simple economic benefit logic, and if all the four shipments can be accommodated in the same truck, then it is way more beneficial for us to send all the four shipments via the same truck, instead of sending four shipments via 4 different trucks. This will require only one truck to travel the same distance to the same customer, however, this time carrying all the four shipments. IN this case, we will save the cost incurred on the three additional trucks (x+x+x) = 3x a, and the total cost we would now be required to bear for transporting all the shipments to the customer is (4x-3x) = x.

                 In the given case, the transportation charges of the 4 trucks are given as:

Truck 1 = 20

Truck 2 = 18

Truck 3 = 15

& Truck 4 = 18

Therefore, the total transport charges = 20 + 18 + 15 + 18 = 71

Now, if we accommodate all four shipments in the same truck, let us say, Truck 1 , then our only cost on transportation will be only 20.

Therefore, we must accept this proposal from the distribution warehouse.